# Hidden Markov Model Filtering with Equal Exit Probabilities

### v1

@@ IV-A Performance of $\alpha-$HMM
Figure 3: The evolution of the instantaneous error probability with iterations under different parameter settings.
IV-B Simulation of Error Probability
In this example, we simulate the error probability of the $\alpha$-HMM algorithm via numerical experiments. We consider a discrete sampling distribution, where the likelihood models of the agents are assumed to belong to a family of binomial distributions with $n=5$ trials and a probability of success given by $p_{\theta}=\sigma\theta$, where $\theta\in\Theta=\{1,2,3\}$. These distributions are characterized by the following probability mass function:
$L(\xi|\theta)=\binom{n}{\xi}p_{\theta}^{\xi}(1-p_{\theta})^{n-\xi},\quad\xi\in\left\{0,1,\ldots,n\right\},$ (60)
where $\binom{n}{\xi}=\frac{n!}{\xi!(n-\xi)!}$. This model setting can be verified to satisfy Assumption 3.
@@ IV-B Simulation of Error Probability
Fig. 3 illustrates the time evolution of the instantaneous error probability defined in (55) under different parameter settings, as estimated through 10,000 Monte Carlo simulations. A clear trend of decay over time is observed, which is consistent with the predictions of Theorem 2. Additionally, when $\sigma$ is fixed and, consequently, $d_{m}$ remains unchanged, increasing $\alpha$ reduces the contraction coefficient $\lambda_{1}$, thereby accelerating the convergence of the error probability. However, it also shifts the fixed point closer to zero, increasing the probability of error occurrence, as described in (56). Conversely, when $\alpha$ is fixed, an increase in $\sigma$ enhances the agent’s ability to distinguish between states, leading to a larger $d_{m}$. This, in turn, results in a smaller contraction coefficient $\lambda_{1}$, further accelerating the convergence of the error probability. Simultaneously, the fixed point moves further away from zero, thereby reducing the probability of error occurrence. These observations further validate the theoretical analysis presented in Theorem 2.
V Conclusion
@@
In this paper, we studied the online optimal filtering problem by assuming that the environmental state follows a Markov chain with equal exit probability and motivated the $\alpha$-HMM algorithm. By conducting a dynamical analysis of the discrete-time stochastic system associated with the algorithm and its deterministic counterpart under steady-state conditions, we investigated the influence of the exit probability $\alpha$ and the agent’s identifiability on the algorithm’s adaptability and learning performance.
@@
Our findings reveal that a larger $\alpha$ enhances convergence speed but lowers confidence in the correct state at the fixed point, increasing the likelihood of erroneous learning. Thus, the choice of $\alpha$ embodies a tradeoff between adaptability and learning precision. Conversely, higher identifiability—representing superior observation quality or a more accurate local model—accelerates convergence, improves adaptability, and shifts the fixed point further from zero, enhancing overall learning performance.

### v2

@@ 5.3 Error Probability of the algorithm
Figure 7: (a) The likelihood function for all states and the true distribution function of the observations under different parameter settings. The dashed lines represent the true distribution functions of the observations, while the solid lines indicate the likelihood functions. (b) The evolution of the instantaneous error probability with iterations under different parameter settings.
@@
Fig. 7(b) illustrates the time evolution of the instantaneous error probability defined in (84) under different parameter settings, which is estimated through 1,000 Monte Carlo simulations. A clear trend of decay over time is observed, consistent with the predictions of Theorem 3.16. Additionally, when $\sigma$ is fixed and $\beta=\alpha$, increasing $\alpha$ reduces the contraction coefficient $\lambda_{1}$, accelerating the convergence of the error probability. However, it also shifts the fixed point closer to zero, increasing the probability of error occurrence, as described in (85). Moreover, regardless of the value of $\sigma$, when $\beta=\alpha$ is sufficiently small, the error probability always converges to zero over time. This observation is consistent with the asymptotic result described in (87). On the other hand, when $\alpha$ and $\beta$ are fixed, an increase in $\sigma$ reduces the filter’s ability to distinguish between states, leading to a smaller $d_{m}$. This results in a larger contraction coefficient $\lambda_{1}$, slowing the convergence of the error probability. Simultaneously, the fixed point moves closer to zero, further increasing the probability of error occurrence. These observations further validate the theoretical analysis presented in Theorem 3.16.
6 Conclusion and Future Work
@@
In this paper, we studied the online optimal filtering problem by assuming that the environmental state follows a Markov chain with equal exit probability and proposed the $\alpha\beta$-HMM algorithm. By conducting a dynamical analysis of the discrete-time stochastic system associated with the algorithm and its deterministic counterpart under steady-state conditions, we investigated the influence of the exit probability $\alpha$, the step-size parameter $\beta$ and the identifiability parameter $\underline{d}$ on the algorithm’s adaptability and learning performance.
@@
Our findings reveal that a larger $\alpha$ enhances convergence speed but lowers confidence in the correct state at the fixed point, increasing the likelihood of erroneous learning. Thus, the choice of $\alpha$ embodies a tradeoff between adaptability and learning precision. When the underlying true state is identifiable, increasing the step-size parameter $\beta$ enhances both learning and adaption performance, but it also amplifies observation noise and should therefore be moderated in high-noise scenarios. Conversely, higher identifiability—representing superior observation quality or a more accurate local model—accelerates convergence, improves adaptability, and shifts the fixed point further from zero, enhancing overall learning performance.
Our analytical approach leverages the contraction properties of nonlinear mappings, proving that the expected discrepancy between the deterministic reference dynamical system and the original stochastic dynamical system decays over time and remains bounded. This allows us to use the fixed point of the deterministic system to provide insights into the algorithm’s instantaneous error probability.
[…]
7 Proof of Lemma 1
For any non-positive input $x\in\left(-\infty,0\right]^{M-1}$, we have $\exp(x_{m})\leq 1$ for $m=1,\ldots,M-1$. Consequently,
$F_{m}(x)$ $=\log\frac{(1-\alpha M)\exp(x_{m})+\alpha+\alpha\sum_{n=1}^{M-1}\exp(x_{n})}{1-\alpha M+\alpha+\alpha\sum_{n=1}^{M-1}\exp(x_{n})}$ (91)
$\leq 0,$
so that $F(x)\in\left(-\infty,0\right]^{M-1}$. For any $x=\left[x_{1},\ldots,x_{M-1}\right]^{\top}$ and $x^{\prime}=\left[x_{1}^{\prime},\ldots,x_{M-1}^{\prime}\right]^{\top}$, by Lagrange’s main value theorem, we have:
$F(x)-F(x^{\prime})=J_{F}(c)\left(x-x^{\prime}\right),$ (92)
where $J_{F}(c)$ is the Jacobian matrix of $F$ at $c=tx+(1-t)x^{\prime}$ with $t\in(0,1)$.
Due to the submultiplicative property of matrix norms:
$\left\|F(x)-F(x^{\prime})\right\|_{\infty}\leq\left\|J_{F}(c)\right\|_{\infty}\left\|x-x^{\prime}\right\|_{\infty}.$ (93)
The function $F_{m}(x)$ is given by:
$F_{m}(x)=\log\Lambda_{m}(x)-\log\Gamma(x),$ (94)
where
$\Lambda_{m}(x)=(1-\alpha M)\exp(x_{m})+\alpha+\alpha\sum_{n=1}^{M-1}\exp(x_{n}),$ (95)
$\Gamma(x)=1-\alpha M+\alpha+\alpha\sum_{n=1}^{M-1}\exp(x_{n}).$ (96)
The partial derivative of $F_{m}(x)$ with respect to $x_{\ell}$ ($\ell\neq m$) is:
$\frac{\partial F_{m}}{\partial x_{\ell}}(x)$ $=\frac{\alpha\exp(x_{\ell})}{\Lambda_{m}(x)}-\frac{\alpha\exp(x_{\ell})}{\Gamma(x)}$ (97)
$=\frac{\alpha(1-\alpha M)\exp(x_{\ell})\left(1-\exp(x_{m})\right)}{\Lambda_{m}(x)\Gamma(x)}.$
Since $0<\alpha<1/M$, $\Lambda_{m}(x)>0$ and $\Gamma(x)>0$, we have:
$\frac{\partial F_{m}}{\partial x_{\ell}}(x)\geq 0,\quad\text{if }x_{m}\leq 0.$ (98)
Similarly, for all $m=1,\ldots,M-1$, we have:
$\frac{\partial F_{m}}{\partial x_{m}}(x)$ $=\frac{(1-\alpha M+\alpha)\exp(x_{m})}{\Lambda_{m}(x)}-\frac{\alpha\exp(x_{m})}{\Gamma(x)}$ (99)
$=\frac{(1-\alpha M)\exp(x_{m})\left(\alpha+\Gamma(x)-\alpha\exp(x_{m})\right)}{\Lambda_{m}(x)\Gamma(x)}.$
Since $\Gamma(x)>\alpha\exp(x_{m})$ for all $m$, we have:
$\frac{\partial F_{m}}{\partial x_{m}}(x)>0.$ (100)
Summing over $l$, the Jacobian satisfies:
$\sum_{\ell=1}^{M-1}\frac{\partial F_{m}}{\partial x_{\ell}}(x)=$ $\frac{(1-\alpha M)\exp(x_{m})+\alpha\sum_{\ell=1}^{M-1}\exp(x_{\ell})}{\Lambda_{m}(x)}$ (101)
$-\frac{\alpha\sum_{\ell=1}^{M-1}\exp(x_{\ell})}{\Gamma(x)}$
$=$ $\frac{\Lambda_{m}(x)-\alpha}{\Lambda_{m}(x)}-\frac{\Gamma(x)-(1-\alpha M+\alpha)}{\Gamma(x)}$
$=$ $\frac{1-\alpha M+\alpha}{\Gamma(x)}-\frac{\alpha}{\Lambda_{m}(x)}.$
For further analysis, by substituting the $\exp(x_{m})$ with $y_{m}$ in the r.h.s. of (101), we can define
$H_{m}(y_{1},\ldots,y_{M-1})\triangleq$ $\frac{1-\alpha M+\alpha}{1-\alpha M+\alpha+\alpha\sum_{n=1}^{m-1}y_{n}}$ (102)
$-\frac{\alpha}{(1-\alpha M)y_{m}+\alpha+\alpha\sum_{n=1}^{M-1}y_{n}},$
where $y_{m}>0$ for all $m=1,\ldots,M-1$. By differentiating with respect to $y_{m}$, it can be verified that:
$\frac{\partial H_{m}}{\partial y_{m}}\geq 0,\quad\text{if }y_{m}\leq 1.$ (103)
Hence when $y_{m}\leq 1$,
$H_{m}(y_{1},\ldots,y_{M-1})$ $\leq H_{m}(y_{1},\ldots,y_{m-1},1,y_{m+1},\ldots,y_{M})$ (104)
$=\frac{1-\alpha M}{1-\alpha M+2\alpha+\sum_{n\neq m}^{M-1}y_{n}}$
$<\frac{1-\alpha M}{1-\alpha M+2\alpha}<1.$
Thus, for $x_{m}\leq 0$, we have $\exp(x_{m})\leq 1$, leading to:
$\sum_{\ell=1}^{M-1}\frac{\partial F_{m}}{\partial x_{\ell}}(x)<\frac{1-\alpha M}{1-\alpha M+2\alpha},\quad\text{if }x_{m}\leq 0.$ (105)
For $x,x^{\prime}\in\left(-\infty,0\right]$, $c=tx+(1-t)x^{\prime}\in\left(-\infty,0\right]$, and for all $m,l=1,\ldots,M-1$, we have:
$\left[J_{F}(c)\right]_{ml}=\frac{\partial F_{m}}{\partial x_{l}}(c)\geq 0,$ (106)
and
$\left\|J_{F}(c)\right\|_{\infty}$ $=\underset{m=1,\ldots,M-1}{\max}\sum_{l=1}^{M-1}\left|\frac{\partial F_{m}}{\partial x_{l}}(c)\right|$ (107)
$=\underset{m=1,\ldots,M-1}{\max}\sum_{l=1}^{M-1}\frac{\partial F_{m}}{\partial x_{l}}(c)$
$<\frac{1-\alpha M}{1-\alpha M+2\alpha}.$
Following (93) we obtain:
$\left\|F(x)-F(x^{\prime})\right\|_{\infty}\leq\frac{1-\alpha M}{1-\alpha M+2\alpha}\left\|x-x^{\prime}\right\|_{\infty},$ (108)
which completes the proof of Lemma 1.
8 Proof of Theorem 2.4
We first show that regardless of the initial value $\hat{x}_{m,0}$, there always exists a time instant $i_{0}$ such that
$\hat{x}_{m,i_{0}}<0,\quad\forall m=1,\ldots,M-1.$ (109)
The recursion $\hat{x}_{m,i}=F_{m}(\hat{x}_{i-1})-\beta d_{m}$ exhibits the following properties:
  • If $\hat{x}_{m,i-1}<0$,
    $\hat{x}_{m,i-1}-\beta d_{m}<\hat{x}_{m,i}<-\beta d_{m},$ (110)
  • If $\hat{x}_{m,i-1}\geq 0$,
    $-\beta d_{m}\leq\hat{x}_{m,i}\leq\hat{x}_{m,i-1}-\beta d_{m}.$ (111)
From (110) and (111), if $\hat{x}_{m,0}\geq 0$, $\hat{x}_{m,i}$ will decrease in each recursion by subtracting $\beta d_{m}$ until it becomes less than 00, and it will remain less than 00 thereafter. Taking
$i_{0}\geq\underset{m=1,\ldots,M-1}{\max}\left\{\frac{\hat{x}_{m,0}}{\beta d_{m}}+1,0\right\},$ (112)
ensures that (109) is satisfied. Additionally, when $\hat{x}_{m,i-1}\geq 0>\hat{x}_{m}^{\infty}$ (which also implies that $\hat{x}_{m,0}\geq 0$), we have:
$\hat{x}_{m,i}\geq-\beta d_{m}\geq\hat{x}_{m}^{\infty}.$ (113)
Thus:
$\left|\hat{x}_{m,i}-\hat{x}_{m}^{\infty}\right|$ $=\hat{x}_{m,i}-\hat{x}_{m}^{\infty}$ (114)
$\leq\hat{x}_{m,i-1}-\beta d_{m}-\hat{x}_{m}^{\infty}$
$=\left(1-\frac{\beta d_{m}}{\hat{x}_{m,i-1}-\hat{x}_{m}^{\infty}}\right)(\hat{x}_{m,i-1}-\hat{x}_{m}^{\infty})$
$\overset{(a)}{\leq}\frac{\hat{x}_{m,0}-\log\frac{\alpha}{1-\alpha M+\alpha}}{\hat{x}_{m,0}-\log\frac{\alpha}{1-\alpha M+\alpha}+\beta d_{m}}(\hat{x}_{m,i-1}-\hat{x}_{m}^{\infty})$
$\leq\frac{\overline{x}_{0}-\log\frac{\alpha}{1-\alpha M+\alpha}}{\overline{x}_{0}-\log\frac{\alpha}{1-\alpha M+\alpha}+\beta\underline{d}}(\hat{x}_{m,i-1}-\hat{x}_{m}^{\infty})$
$=\frac{\overline{x}_{0}-\log\frac{\alpha}{1-\alpha M+\alpha}}{\overline{x}_{0}-\log\frac{\alpha}{1-\alpha M+\alpha}+\beta\underline{d}}\left|\hat{x}_{m,i-1}-\hat{x}_{m}^{\infty}\right|.$
Here, step $(a)$ holds since
$\hat{x}_{m,i}\leq\hat{x}_{m,0}-i\beta d_{m}\leq\hat{x}_{m,0},$ (115)
and by Lemma 2.2:
$\hat{x}_{m}^{\infty}$ $\geq\log\frac{\alpha}{\alpha\exp(\beta d_{m})+(1-\alpha M)(\exp(\beta d_{m})-1)}$ (116)
$\geq\log\frac{\alpha}{(1-\alpha M+\alpha)\exp(\beta d_{m})}$
$=\log\frac{\alpha}{1-\alpha M+\alpha}-\beta d_{m}$
Next, we invoke the following lemma to complete the proof:
Lemma 8.19 (Banach Fixed Point Theorem[31]).
If $T:X\to X$ is a contraction on a complete metric space $(X,\rho)$, then there is a unique fixed point $x\in X$ (35).
Starting from $i_{0}$, the sequence $\hat{x}_{i}$ will always remain within $\left(-\infty,0\right]^{M-1}$. Therefore, by Lemma 1 and Lemma 8.19, there exists a unique fixed point $\hat{x}^{\infty}\in\left(-\infty,0\right]^{M-1}$, such that
$\lim_{i\to\infty}\hat{x}_{i}=\hat{x}^{\infty}.$ (117)
From the proof of Lemma 1, when $\hat{x}_{m,i-1}\leq 0$:
$\left|\hat{x}_{m,i}-\hat{x}^{\infty}_{m}\right|$ $=\left|F_{m}(\hat{x}_{i-1})-F_{m}(\hat{x}^{\infty})\right|$ (118)
$=\left|\sum_{\ell=1}^{M-1}\left[J_{F}(c_{i-1})\right]_{m\ell}\left(\hat{x}_{\ell,i-1}-\hat{x}_{\ell}^{\infty}\right)\right|$
$\overset{(b)}{\leq}\left(\sum_{\ell=1}^{M-1}\left[J_{F}(c_{i-1})\right]_{m\ell}\right)\left\|\hat{x}_{i-1}-\hat{x}^{\infty}\right\|_{\infty}$
(105)${\leq}\frac{1-\alpha M}{1-\alpha M+2\alpha}\left\|\hat{x}_{i-1}-\hat{x}^{\infty}\right\|_{\infty}.$
Here $c_{i-1}=t\hat{x}_{i-1}+(1-t)\hat{x}^{\infty}$ with $t\in(0,1)$, and step $(b)$ holds since $c_{m,i-1}\leq 0$ and from the proof of Lemma 1,
$\left[J_{F}(c_{i-1})\right]_{m\ell}\geq 0,\quad 1\leq m,\ell\leq M-1.$ (119)
Combining (114) and (118) together, we conclude that (42) holds.
9 Proof of Theorem 2.11
From Lemma 2.2, for all $m=1,\ldots,M-1$, we have
$\exp(\hat{x}_{m}^{\infty})=\frac{\alpha}{\alpha\exp(\beta d_{m})+(1-\alpha M)(\exp(\beta d_{m})-1)\hat{\mu}_{0}^{\infty}},$ (120)
which leads to
$\alpha\exp(\beta d_{m})\exp(\hat{x}_{m}^{\infty})$ (121)
$+(\exp(\beta d_{m})-1)(1-\alpha M)\hat{\mu}_{0}^{\infty}\exp(\hat{x}_{m}^{\infty})=\alpha.$
Summing over $m=1,\ldots,M-1$, we have:
$\alpha(M-1)$ (122)
$=$ $\left((1-\alpha M)\hat{\mu}_{0}^{\infty}+\alpha\right)\sum_{m=1}^{M-1}\exp(\hat{x}_{m}^{\infty})\exp(\beta d_{m})$
$-(1-\alpha M)\hat{\mu}_{0}^{\infty}\sum_{m=1}^{M-1}\exp(\hat{x}_{m}^{\infty}).$
Using (73), $\sum_{m=1}^{M-1}\exp(\hat{x}_{m}^{\infty})=1/\hat{\mu}_{0}^{\infty}-1$, we obtain:
$1-\alpha-(1-\alpha M)\hat{\mu}_{0}^{\infty}$ (123)
$=$ $\left((1-\alpha M)\hat{\mu}_{0}^{\infty}+\alpha\right)\sum_{m=1}^{M-1}\exp(\beta d_{m})\exp(\hat{x}_{m}^{\infty})$
$\geq$ $\left((1-\alpha M)\hat{\mu}_{0}^{\infty}+\alpha\right)\exp(\beta\underline{d})(1/\hat{\mu}_{0}^{\infty}-1).$
Since $\hat{\mu}_{0}^{\infty}<1$, we have
$1-\alpha>(1-\alpha M)\exp(\beta\underline{d})-(1-\alpha M)(\exp(\beta\underline{d})-1)\hat{\mu}_{0}^{\infty}.$ (124)
With $01-\alpha M>$ 0 and $\exp(\underline{d})>1$ from Assumption 2, this inequality establishes the lower bound:
$\hat{\mu}_{0}^{\infty}>\frac{(1-\alpha M)\exp(\beta\underline{d})-1+\alpha}{(1-\alpha M)(\exp(\beta\underline{d})-1)}=\underline{\mu}.$ (125)
On the other hand, we also have
$1-\alpha-(1-\alpha M)\hat{\mu}_{0}^{\infty}$ (126)
$\leq$ $\left((1-\alpha M)\hat{\mu}_{0}^{\infty}+\alpha\right)\exp(\beta\overline{d})(1/\hat{\mu}_{0}^{\infty}-1),$
Using the previously derived lower bound for $\hat{\mu}_{0}^{\infty}$, the upper bound can be similarly obtained:
$\hat{\mu}_{0}^{\infty}<\frac{(1-\alpha M-\alpha+\alpha/\underline{\mu})\exp(\beta\overline{d})-1+\alpha}{(1-\alpha M)(\exp(\beta\overline{d})-1)}.$ (127)
This completes the proof.
10 Proof of Lemma 3.14
We first show that, under Assumption 3 and the dynamics in (80), the log-belief ratio can be bounded almost surely.
Lemma 10.20.
The log-belief ratio $\boldsymbol{x}_{m,i}$ in (80) satisfies:
$q_{m}\leq\boldsymbol{x}_{m,i}\leq Q_{m}\quad a.s.,$ (128)
where
$q_{m}=\log\frac{\alpha}{1-\alpha M+\alpha}-\beta d_{m}-\beta C,$ (129)
$Q_{m}=\log\frac{1-\alpha M+\alpha}{\alpha}-\beta d_{m}+\beta C.$ (130)
$C$ is defined as in (78).
Proof 10.21.
Supplementary Material III.
Define $\boldsymbol{\delta}_{m,i}=\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})}+d_{m}$. Then, the original and reference dynamics can be expressed as:
$\boldsymbol{x}_{m,i}=F_{m}(\boldsymbol{x}_{1,i-1},\ldots,\boldsymbol{x}_{M-1,i-1})-\beta d_{m}+\beta\boldsymbol{\delta}_{m,i},$ (131)
and
$\hat{x}_{m,i}=F_{m}(\hat{x}_{1,i-1},\ldots,\hat{x}_{M-1,i-1})-\beta d_{m}.$ (132)
For any $x$, if $0\leq x_{m}\leq Q_{m}$, notice that $F_{m}(\hat{x}^{\infty})-\beta d_{m}=\hat{x}_{m}^{\infty}$, we have:
$\left|F_{m}(x)-F_{m}(\hat{x}^{\infty})\right|$ (133)
$=$ $F_{m}(x)-F_{m}(\hat{x}^{\infty})$
$\leq$ $x_{m}-\beta d_{m}-\hat{x}_{m}^{\infty}$
$\leq$ $\left(1-\frac{\beta d_{m}}{x_{m}-\hat{x}_{m}^{\infty}}\right)(x_{m}-\hat{x}_{m}^{\infty})$
$\overset{(a)}{\leq}$ $\left(1-\frac{\beta d_{m}}{2\log\frac{1-\alpha M+\alpha}{\alpha}+\beta C}\right)(x_{m}-\hat{x}_{m}^{\infty})$
$\leq$ $\left(1-\frac{\beta\underline{d}}{2\log\frac{1-\alpha M+\alpha}{\alpha}+\beta C}\right)\left|x_{m}-\hat{x}_{m}^{\infty}\right|,$
here step $(a)$ follows from Lemma 10.20 and (116).
If $q_{m}\leq x_{m}<0$, then from Lemma 1:
$\left|F_{m}(x)-F_{m}(\hat{x}^{\infty})\right|=$ $\left|\sum_{\ell=1}^{M-1}\left[J_{F}(c)\right]_{m\ell}\left(x_{\ell}-\hat{x}_{\ell}^{\infty}\right)\right|$ (134)
$\overset{(a)}{\leq}$ $\left(\sum_{\ell=1}^{M-1}\left[J_{F}(c)\right]_{m\ell}\right)\left\|x-\hat{x}^{\infty}\right\|_{\infty}$
(${\leq}$ $\frac{1-\alpha M}{1-\alpha M+2\alpha}\left\|x-\hat{x}^{\infty}\right\|_{\infty}.$
Here $c=tx+(1-t)\hat{x}^{\infty}$ with $t\in(0,1)$, and step $(a)$ holds since $c_{m}<0$ and from the proof of Lemma 1,
$\left[J_{F}(c)\right]_{m\ell}>0,\quad 1\leq m,\ell\leq M-1.$ (135)
Let $Q=\underset{m=1,\ldots,M-1}{\max}\left\{\left|Q_{m}\right|,\left|q_{m}\right|\right\}$, for all $x\in\left[-Q,Q\right]^{M-1}$ we have:
$\left\|F(x)-F(\hat{x}^{\infty})\right\|_{\infty}\leq\lambda_{1}\left\|x-\hat{x}^{\infty}\right\|_{\infty},$ (136)
where $\lambda_{1}$ is as defined in (83).
From the recursion (131) and (132), we have:
$\mathbb{E}\left[\left.\left\|\boldsymbol{x}_{i}-\hat{x}^{\infty}\right\|_{\infty}\right|\boldsymbol{x}_{i-1}\right]$ (137)
$=$ $\mathbb{E}\left[\left.\left\|F\left(\boldsymbol{x}_{i-1}\right)-F\left(\hat{x}^{\infty}\right)+\beta\boldsymbol{\delta}_{i}\right\|_{\infty}\right|\boldsymbol{x}_{i-1}\right]$
$\leq$ $\mathbb{E}\left[\left.\left\|F\left(\boldsymbol{x}_{i-1}\right)-F\left(\hat{x}^{\infty}\right)\right\|_{\infty}\right|\boldsymbol{x}_{i-1}\right]+\beta\mathbb{E}\left[\left\|\boldsymbol{\delta}_{i}\right\|_{\infty}\right]$
$\overset{(b)}{\leq}$ $\lambda_{1}\mathbb{E}\left[\left.\left\|\boldsymbol{x}_{i-1}-\hat{x}^{\infty}\right\|_{\infty}\right|\boldsymbol{x}_{i-1}\right]+\beta C.$
Here step $(b)$ is due to (136) along with Assumption 3.
By taking expectation again, we obtain
$\mathbb{E}\left[\left\|\boldsymbol{x}_{i}-\hat{x}^{\infty}\right\|_{\infty}\right]$ (138)
$\leq$ $\lambda_{1}\mathbb{E}\left[\left\|\boldsymbol{x}_{i-1}-\hat{x}^{\infty}\right\|_{\infty}\right]+\beta C$
$\leq$ $\lambda_{1}^{i}\mathbb{E}\left[\left\|\boldsymbol{x}_{0}-\hat{x}^{\infty}\right\|_{\infty}\right]+\frac{\beta C(1-\lambda_{1}^{i})}{1-\lambda_{1}}$
$\leq$ $O(\lambda_{1}^{i})+\frac{\beta C}{1-\lambda_{1}}.$
11 Proof of Theorem 3.16
The instantaneous error probability can be computed as:
$p_{i}^{e}=$ $\mathbb{P}\left[\;\exists\;\theta_{m}\neq\theta_{0},\text{ s.t. }\boldsymbol{\mu}_{i}(\theta_{m})\geq\boldsymbol{\mu}_{i}(\theta_{0})\right]$ (139)
$=$ $\mathbb{P}\left[\;\exists\;\theta_{m}\neq\theta_{0},\text{ s.t. }\boldsymbol{x}_{m,i}\geq 0\right]$
$\leq$ $\mathbb{P}\left[\;\exists\;\theta_{m}\neq\theta_{0},\text{ s.t. }\boldsymbol{x}_{m,i}-\hat{x}_{m}^{\infty}\geq-\overline{x}^{\infty}\right]$
$\leq$ $\mathbb{P}\left[\underset{m=1,\ldots,M-1}{\max}\left|\boldsymbol{x}_{m,i}-\hat{x}_{m}^{\infty}\right|\geq-\overline{x}^{\infty}\right]$
$=$ $\mathbb{P}\left[\left\|\boldsymbol{x}_{i}-\hat{x}^{\infty}\right\|_{\infty}\geq-\overline{x}^{\infty}\right]$
$\overset{(a)}{\leq}$ $\frac{\mathbb{E}\left[\left\|\boldsymbol{x}_{i}-\hat{x}^{\infty}\right\|_{\infty}\right]}{-\overline{x}^{\infty}}$
$\overset{(b)}{\leq}$ $\frac{O(\lambda_{1}^{i})+\beta C}{-(1-\lambda_{1})\overline{x}^{\infty}}.$
where step $(a)$ is due to Markov’s inequality and step $(b)$ follows from Lemma 3.14.
Furthermore, due to the fact that:
$\lambda_{1}<1,$ (140)
we always have:
$\underset{i\to\infty}{\lim\sup}\;p_{i}^{e}\leq\frac{\beta C}{-(1-\lambda_{1})\overline{x}^{\infty}}.$ (141)
If $\alpha\to 0$, we have:
$\frac{\beta\underline{d}}{2\log\frac{1-\alpha M+\alpha}{\alpha}+\beta C}=O(\frac{1}{\log\alpha}),$ (142)
thus
$1-\lambda_{1}=\min\left\{\frac{2\alpha}{1-\alpha M+2\alpha},\frac{\beta\underline{d}}{2\log\frac{1-\alpha M+\alpha}{\alpha}+\beta C}\right\}=O(\alpha).$ (143)
Moreover, from (74) and (37), we have:
$\lim_{\alpha\to 0}\hat{\mu}_{0}^{\infty}=1,$ (144)
and then for all $m=1,\ldots,M-1$,
$\lim_{\alpha\to 0}\hat{x}_{m}^{\infty}=-\infty.$ (145)
If $\beta=O(\alpha)$, then:
$\lim_{\alpha\to 0}\left(\underset{i\to\infty}{\lim\sup}\;p_{i}^{e}\right)\leq\frac{\beta C}{-(1-\lambda_{1})\overline{x}^{\infty}}=\frac{O(1)}{-\overline{x}^{\infty}}=0.$ (146)
On the other hand, as $p_{i}^{e}$ is a probability measure, it is obvious that:
$\lim_{\alpha\to 0}\left(\underset{i\to\infty}{\lim\sup}\;p_{i}^{e}\right)\geq 0.$ (147)
By applying the squeeze theorem for limits, we then have:
$\lim_{\alpha\to 0}\left(\limsup_{i\to\infty}p_{i}^{e}\right)=0.$ (148)